The number density of free electrons in a copper conductor estimated to be 8.5 x 10^28 /m³. How long does an electron take, to drift from one end of a wire 3.0 m long to its other end? The area of cross section of the copper wire is 2.0 x 10^-6 m² and it is carrying a current of 3.0 A

Grade:12

1 Answers

Ruhi
6 years ago
If V.d is drift velocity and I is current V.d=I÷enA .Then 3÷{1.6×10^-19×2×10^-6×8.5×10-28}=1.102×10^-4Or 0.000102